OpenStudy (anonymous):

What number is halfway between 1 and 5/6?

jimthompson5910 (jim_thompson5910):

hint: add the two numbers up, then divide that sum by 2

OpenStudy (anonymous):

Umm okay thank you!

jimthompson5910 (jim_thompson5910):

you're welcome, tell me what you get

OpenStudy (anonymous):

I got 4/7 but I'm not sure if I did it right.

jimthompson5910 (jim_thompson5910):

that is incorrect

OpenStudy (anonymous):

Umm okay thank you

jimthompson5910 (jim_thompson5910):

can you show me how you got 4/7?

OpenStudy (anonymous):

I did what you said to do by taking the numbers and adding them together, after I added the numbers together I got 4/7. But I probably did some math wrong

OpenStudy (jdoe0001):

hmm, getting "7" as lcd shows something ... went really awry or ... is made up hehe

jimthompson5910 (jim_thompson5910):

the number 1 can be thought of as 6/6 since 6/6 = 1 this means that 1 + 5/6 is really the same as 6/6 + 5/6 with me so far?

OpenStudy (anonymous):

Okay so then would I add the 5/6 and 6/6 together?

jimthompson5910 (jim_thompson5910):

correct, how would you do that?

OpenStudy (anonymous):

You would only add the 5 and 6 together, right?

jimthompson5910 (jim_thompson5910):

correct, you add the numerators and place that over the common denominator of 6

OpenStudy (anonymous):

Okay so would the final answer be 11/6?

jimthompson5910 (jim_thompson5910):

now cut that in half or multiply that by 1/2

jimthompson5910 (jim_thompson5910):

11/6 * 1/2 = ???

OpenStudy (anonymous):

Would the denominator would be three?

jimthompson5910 (jim_thompson5910):

no, you don't divide 6 by 2 to get 3

OpenStudy (anonymous):

Okay.. So do I divide the whole fraction by a 1/2 somehow?

jimthompson5910 (jim_thompson5910):

to multiply 11/6 by 1/2, you multiply the numerators together and the denominators together multiply numerators: 11*1 = 11 multiply denominators: 6*2 = 12 So 11/6 * 1/2 = 11/12

jimthompson5910 (jim_thompson5910):

that means the number 11/12 is halfway between 1 and 5/6

OpenStudy (anonymous):

Oh okay thank you for helping and explaining it! It helped me understand it better to where I can do all the other ones myself! Thank you again(:

jimthompson5910 (jim_thompson5910):

glad to be of help, you're welcome

OpenStudy (goformit100):

First of all @justgottabehappy98 : A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct